# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

### 27 Feb 2024 v1

@@ 4 Spectral distribution of the PESS preconditioned matrix
$2\Re(\mu)$$\geq$ $\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)}{\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}\frac{v^{*}B\Lambda_{1}^{-1}B^{T}v}{v^{*}\Lambda_{2}v}$ (4.42)
$\geq$ $\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)\lambda_{\min}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}.$
Hence, the proof of part $(2).$ follows by merging the inequalities of (4.35) and (4.42). $\blacksquare$ Utilizing the established bounds in Theorem 4.3, we derive the subsequent estimations for non-real eigenvalues of SS and EGSS preconditioned matrices.
@@
Let ${\mathscr{P}}_{SS}$ be defined as in (1.16) with $\alpha>0$ and let $\lambda$ be a non-real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then, $(1).$   If $Cv=0,$ then $\lambda$ satisfies
$\frac{2\lambda_{\min}(A)}{2\alpha+\lambda_{\min}(A)}\leq|\lambda|\leq\sqrt{\frac{4\lambda_{\max}(BB^{T})}{\alpha^{2}+\alpha\lambda_{\min}(A)+\lambda_{\max}(BB^{T})}}.$
$(2).$    If $Cv\neq 0,$ the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
[…]
$\left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\frac{\sqrt{\lambda_{\max}(BB^{T})+\lambda_{\max}(C^{T}C)}}{\alpha}.$
@@
Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.17), and let $P,Q$ and $W$ are SPD matrices with $\alpha,\beta,\gamma>0.$ Let $\lambda$ be a non-real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then $(1).$ If $Cv=0,$ then $\lambda$ satisfies
$\frac{2\lambda_{\min}(P^{-1}A)}{2\alpha+\lambda_{\min}(P^{-1}A)}\leq|\lambda|\leq\sqrt{\frac{4\lambda_{\max}(Q^{-1}BP^{-1}B^{T})}{\alpha\beta+\beta\lambda_{\min}(P^{-1}A)+\lambda_{\max}(Q^{-1}BP^{-1}B^{T})}}.$
$(2).$   If $Cv\neq 0,$ then the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
$\frac{\lambda_{\min}(P^{-1}A)\,\lambda_{\min}(Q^{-1}BP^{-1}B^{T})}{2\left(\beta\lambda_{\max}^{2}(P^{-1}A)+\alpha\lambda_{\max}(Q^{-1}BP^{-1}B^{T})+\frac{\alpha^{2}}{\gamma}\lambda_{\max}(Q^{-1}C^{T}W^{-1}C)\right)}$
[…]
$\left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\sqrt{\frac{1}{\alpha\beta}\lambda_{\max}(Q^{-1}BP^{-1}B^{T})+\frac{1}{\beta\gamma}\lambda_{\max}(\Lambda_{2}^{-1}C^{T}W^{-1}C)}.$
@@ 5 Numerical experiments
In this section, we conduct a few numerical experiments to showcase the superiority and efficiency of the proposed PESS preconditioner over the existing preconditioners to enhance the convergence speed of the Krylov subspace iterative process to solve three-by-three block SPPs. Our study involves a comparative analysis among GMRES process and the PGMRES process employing the proposed PESS preconditioner ${\mathscr{P}}_{PESS}$ and the existing preconditioners ${\mathscr{P}}_{BD}$, ${\mathscr{P}}_{IBD}$, ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS}.$ The numerical results are reported from the aspect of iteration counts (abbreviated as “IT”) and elapsed CPU times in seconds (abbreviated as “CPU”). Each subsystem featured in Algorithm 1 is precisely solved by applying the Cholesky factorization of the coefficient matrices. For all iterative process, the initial guess vector is ${\bf u}_{0}=0\in{\mathbb{R}}^{n+m+p}$, and the termination criterion is
@@
SS IT $10$ $9$ $9$ $9$ $9$
CPU $0.2879$ $2.0186$ $12.7565$ $63.8265$ $211.2261$
RES $071.4343e-$ 07 $077.5194e-$ 07 $076.0005e-$ 07 $075.2050e-$ 07 $074.6709e-$ 07
RSS IT $9$ $9$ $9$ $9$ $9$
CPU $0.1990$ $1.8589$ $12.7943$ $62.9284$ $211.9332$
RES $077.8940e-$ 07 $075.4297e-$ 07 $074.5044e-$ 07 $073.8357e-$ 07 $073.3306e-$ 07
EGSS IT $9$ $9$ $9$ $9$ $9$
CPU $0.2130$ $1.4008$ $8.7647$ $41.7527$ $156.1426$
RES $071.1981e-$ 07 $071.7824e-$ 07 $071.5671e-$ 07 $071.4056e-$ 07 $071.2878e-$ 07
${}^{\dagger}$ IT $3$ $3$ $3$ $3$ $3$
CPU $0.2344$ $0.9795$ 5.8979 28.3526 98.0878
$s=12$ RES $081.0992e-$ 08 $082.4637e-$ 08 $084.8787e-$ 08 $087.2624e-$ 08 $088.4769e-$ 08
${}^{\dagger}$ IT $3$ $3$ $3$ $3$ $3$
CPU $0.1929$ $0.9965$ $5.9099$ $28.0443$ $101.0273$
$s=13$ RES $098.6456e-$ 09 $081.9393e-$ 08 $083.8703e-$ 08 $085.8810e-$ 08 $087.0253e-$ 08
@@
Example 1.
We consider the three-by-three block SPP (1.1) taken from [18], where
$A=\left[\begin{matrix}I\otimes G+G\otimes I&0\\ 0&I\otimes G+G\otimes I\end{matrix}\right]\in{\mathbb{R}}^{2l^{2}\times 2l^{2}},\,\,B=\left[\begin{matrix}I\otimes F&F\otimes I\end{matrix}\right]\in{\mathbb{R}}^{l^{2}\times 2l^{2}},$ and $C=E\otimes F\in{\mathbb{R}}^{l^{2}\times l^{2}},$ where $G=\frac{1}{(l+1)^{2}}\,\mathrm{tridiag}(-1,2,-1)\in{\mathbb{R}}^{l\times l},\quad F=\frac{1}{l+1}\,\mathrm{tridiag}(0,1,-1)\in{\mathbb{R}}^{l\times l}$ , and $E=\mathrm{diag}(1,l+1,\ldots,l^{2}-l+1)\in{\mathbb{R}}^{l\times l}.$ Here, $E\otimes F$ denotes the Kronecker product of two matrices $E$ and $F,$ and $\mathrm{tridiag}(a,b,c)\in{\mathbb{R}}^{l\times l}$ denotes the $l\times l$ tridiagonal matrix with diagonal, subdiagonal, and superdiagonal entries are $a,b,$ and $c,$ respectively. For this problem, the size of the matrix $\mathcal{A}$ is $4l^{2}.$ We consider the parameter choices for the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},{\mathscr{P}}_{EGSS}$ and ${\mathscr{P}}_{PESS}$ in two cases.
  • In Case I: $\alpha=0.2$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.2,\beta=2,\gamma=0.002$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha I,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma I$ for ${\mathscr{P}}_{PESS}.$
  • In Case II: $\alpha=2$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=2,\beta=2,\gamma=0.002$ and $P=I,Q=I,W=CC^{T}$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha I,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma CC^{T}$ for ${\mathscr{P}}_{PESS}.$
The optimal values (experimentally) for $s$ in the range $[10,20]$ are determined to be $12$ and $13$ for minimal CPU times. Following [18], selection of $\hat{A}$ and $\hat{S}$ (SPD approximations of $A$ and $S,$ respectively) for IBD preconditioner are done as follows:
$\hat{A}=LL^{T},\quad\hat{S}=\mathrm{diag}(B\hat{A}^{-1}B^{T}),$
@@
Figure 1: Convergence curves for IT versus RES of PGMRES process employing BD, IBD, SS, RSS, EGSS, and PESS $(s=13)$ preconditioners in Case II for Example 1.
@@
(g) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=12$
Figure 2: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
@@
Figure 1 illustrates convergence curves pertaining to preconditioners BD, IBD, SS, RSS, EGSS, and PESS ($s=13$) for Case II with $l=16,32,48,64$ and $80$ These curves depict the relationship between the relative residue at each iteration step (RES) and IT counts. To further illustrate the superiority of the PESS preconditioner, spectral distributions of $\mathcal{A}$ and the preconditioned matrices ${\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{RSS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for the Case II with $l=16$ and $s=12$ are displayed in Figure 2. To investigate the robustness of the proposed PESS preconditioner, we measure the condition number of the ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ which is for any nonsingular matrix $A$ defined by $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2}.$ In Figure 3, the influence of the parameter $s$ on the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ is depicted. The parameter $s$ is chosen from the interval $[5,50]$ with step size one. The data presented in Table 2 and Figures 1-3 allow us to make the following noteworthy observations:
@@
  • From Table 2, it is observed that the GMRES has a very slow convergence speed. In both Cases I and II, our proposed PESS preconditioner outperforms all other compared existing preconditioners from the aspects of IT and CPU times. For example, in Case II with $l=80,$ our proposed PESS preconditioner is $71\%$, $30\%$, $54\%$, $54\%$, and $37\%$ more efficient than the existing BD, IBD, SS, RSS, and EGSS preconditioner, respectively. Similar patterns are noticed for $l=16,32,48$ and $64$.
@@
  • In Figure 1, it is evident that the PESS preconditioner has a faster convergence speed than the other baseline preconditioners when applied to the PGMRES process for all $l=16,32,48,64$ and $80$

### 17 May 2024 v2

@@ 2. $(2)$
$2\Re(\mu)$$\geq$ $\frac{\xi_{\min}}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}\frac{v^{*}B\Lambda_{1}^{-1}B^{T}v}{v^{*}\Lambda_{2}v}\geq\frac{\xi_{\min}\eta_{\min}}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}.$ (4.44)
Hence, the proof of part $(2)$ follows by merging the inequalities of (4.37) and (4.44). $\blacksquare$
Remark 4.3.
From the bounds in $(2)$ of Theorem 4.3, we obtain
${\frac{1}{|\tau|^{2}}\left(\frac{\xi_{\min}\eta_{\min}}{2(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max})}+\frac{1}{s}\right)\leq\frac{1}{s}-\Re(\lambda)\leq\frac{1}{|\tau|^{2}}(\frac{\xi_{\max}}{2}+\frac{1}{s})}$
and
$|\Im(\lambda)|\leq\frac{1}{s\tau\bar{\tau}}\sqrt{\eta_{\max}+\theta_{\max}},$
where $\tau=s\mu+1.$ Thus, the real and imaginary parts of the eigenvalues of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ cluster better as $s$ grows. Therefore, the PESS preconditioner can accelerate the rate of convergence of the Krylov subspace process, like $GMRES$.
Utilizing the established bounds in Theorem 4.3, we derive the subsequent estimations for non-real eigenvalues of SS and EGSS preconditioned matrices.
@@
Let ${\mathscr{P}}_{SS}$ be defined as in (1.10) with $\alpha>0$ and let $\lambda$ be a non-real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then,
$(1)$   If $Cv=0,$ $\lambda$ satisfies
$\frac{2\kappa_{\min}}{2\alpha+\kappa_{\min}}\leq|\lambda|\leq\sqrt{\frac{4\tau_{\max}}{\alpha^{2}+\alpha\kappa_{\min}+\tau_{\max}}}.$
$(2)$    If $Cv\neq 0,$ the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
$\frac{\kappa_{\min}\,\tau_{\min}}{2\alpha\left(\kappa^{2}_{\max}+\tau_{\max}+\beta_{\max}\right)}\leq\Re\left(\frac{\lambda}{2-\lambda}\right)\leq\frac{\kappa_{\max}}{2\alpha}\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\frac{\sqrt{\tau_{\max}+\beta_{\max}}}{\alpha},$
where $\kappa_{\min}=\lambda_{\min}(A),$ $\tau_{\max}=\lambda_{\max}(BB^{T}),$ $\tau_{\min}=\lambda_{\min}(BB^{T})$ and $\beta_{\max}=\lambda_{\max}(C^{T}C).$
@@
Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.11), and let $P,Q$ and $W$ are SPD matrices with $\alpha,\beta,\gamma>0.$ Let $\lambda$ be a non-real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then $(1)$ If $Cv=0,$ $\lambda$ satisfies
$\frac{2\tilde{\kappa}_{\min}}{2\alpha+\tilde{\kappa}_{\min}}\leq|\lambda|\leq\sqrt{\frac{4\tilde{\tau}_{\max}}{\alpha\beta+\beta\tilde{\kappa}_{\min}+\tilde{\tau}_{\max}}}.$
$(2)$   If $Cv\neq 0,$ the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
$\frac{\tilde{\kappa}_{\min}\,\tilde{\tau}_{\min}}{2\left(\beta\tilde{\kappa}^{2}_{\max}+\alpha\tilde{\tau}_{\max}+(\alpha^{2}/\gamma)\tilde{\beta}_{\max}\right)}\leq\Re\left(\frac{\lambda}{2-\lambda}\right)\leq\frac{\tilde{\kappa}_{\max}}{2\alpha}\leavevmode\nobreak\ \mbox{and}$
$\left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\sqrt{\frac{1}{\alpha\beta}\tilde{\tau}_{\max}+\frac{1}{\beta\gamma}\tilde{\beta}_{\max}},$
where $\tilde{\kappa}_{\min}=\lambda_{\min}(P^{-1}A),$ $\tilde{\tau}_{\max}=\lambda_{\max}(Q^{-1}BP^{-1}B^{T}),$ $\tilde{\tau}_{\min}=\lambda_{\min}(Q^{-1}BP^{-1}B^{T})$ and $\tilde{\beta}_{\max}=\lambda_{\max}(C^{T}C).$
@@
5 Local PESS (LPESS) preconditioner
To enhance the efficiency of the PESS preconditioner, in this section, we propose a relaxed version of the PESS preconditioner by incorporating the concept of RSS preconditioner [9]. By omitting the term $\Lambda_{1}$ from the $(1,1)$-block of ${\mathscr{P}}_{PESS}$, we present the local PESS (LPESS) preconditioner, denoted as ${\mathscr{P}}_{LPESS}$, defined as follows:
${\mathscr{P}}_{LPESS}:=\left[\begin{matrix}A&sB^{T}&0\\ -sB&\Lambda_{2}&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right].$ (5.1)
The implementation of the LPESS preconditioner is similar to the Algorithm 1. However, there is one modification in step $3,$ i.e., we need to solve a linear subsystem $(sA+s^{2}B^{T}\widehat{X}^{-1}B)w_{1}=v$ instead of $\widehat{A}w_{1}=v.$
To illustrate the efficiency of the LPESS preconditioner, we study the spectral distribution of the preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}.$ To achieve this, we consider the following decomposition of the matrices $\mathcal{A}$ and ${\mathscr{P}}_{LPESS}$ as follows:
$\mathcal{A}=\mathfrak{L}\mathfrak{D}\mathfrak{U}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ {\mathscr{P}}_{LPESS}=\mathfrak{L}\widetilde{\mathfrak{D}}\mathfrak{U},$ (5.2)
where
$\mathfrak{L}=\left[\begin{matrix}I&0&0\\ -BA^{-1}&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{U}=\left[\begin{matrix}I&A^{-1}B^{T}&0\\ 0&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{D}=\left[\begin{matrix}A&0&0\\ 0&Q&-C^{T}\\ 0&C&0\end{matrix}\right],\widetilde{\mathfrak{D}}=\left[\begin{matrix}sA&0&0\\ 0&\Lambda_{2}+sQ&-C^{T}\\ 0&C&\Lambda_{3}\end{matrix}\right],$
and $Q=BA^{-1}B^{T}.$ Using the decomposition in (5.2), we have
${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}=\mathfrak{U}^{-1}\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right]\mathfrak{U},$ (5.3)
where $M=\left[\begin{matrix}\Lambda_{2}+sQ&-sC^{T}\\ sC&\Lambda_{3}\end{matrix}\right]$ and $K=\left[\begin{matrix}Q&-C^{T}\\ C&0\end{matrix}\right].$ Since, ${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}$ is similar to $\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right],$ ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has eigenvalue $\frac{1}{s}$ with multiplicity $n$ and the remaining eigenvalues satisfies the generalized eigenvalue problem $Kp=\lambda Mp.$
Theorem 5.1.
Let $A\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ and $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Assume that $s>0,$ then LPESS preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has $n$ eigenvalue equal to $\frac{1}{s}.$ The remaining $m+p$ eigenvalues satisfies the following:
  1. $(1)$
    real eigenvalues located in the interval
    $\Big{[}\min\Big{\{}\frac{\vartheta_{\min}}{1+s\vartheta_{\min}},\frac{\theta_{\min}}{\vartheta_{\max}+s\theta_{\min}}\Big{\}},\frac{\vartheta_{\max}}{1+s\vartheta_{\max}}\Big{]}.$
  2. $(2)$
    If $\lambda$ is any non-real eigenvalue (i.e., $\Im(\lambda)\neq 0$), then
    $\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\leq|\lambda|\leq\sqrt{\frac{\theta_{\max}}{1+s\vartheta_{\min}+s^{2}\theta_{\max}}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \frac{1}{s(1+s\sqrt{\theta_{\max}})}\leq|\lambda-\frac{1}{s}|\leq\frac{2}{s(2+s\vartheta_{\min})},$
where $\vartheta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}Q),$ $\vartheta_{\min}:=\lambda_{\min}(\Lambda_{2}^{-1}Q)$ and $\theta_{\min}:=\lambda_{\min}(\Lambda_{3}^{-1}C\Lambda_{1}^{-1}C^{T}).$
Proof.
Observe that,
$M^{-1}K=\mathcal{F}^{-1}\widetilde{M}^{-1}\widetilde{K}\mathcal{F},$ (5.4)
where $\mathcal{F}=\left[\begin{matrix}I&0\\ 0&-I\end{matrix}\right],$ $\widetilde{M}=\left[\begin{matrix}\Lambda_{2}+sQ&C^{T}\\ -C&\Lambda_{3}\end{matrix}\right]$ and $\widetilde{K}=\left[\begin{matrix}Q&C^{T}\\ -C&0\end{matrix}\right].$ Hence, $M^{-1}K$ is similar to $\widetilde{M}^{-1}\widetilde{K}.$ Now, $\widetilde{M}^{-1}\widetilde{K}$ is in the form of the preconditioned matrix discussed in [28]. Therefore, by applying the Corollary 4.1 and Theorem 4.3 of [28], we obtain the desired bounds in $(1)$ and $(2).$blacksquare
@@ 6 Numerical experiments
In this section, we conduct a few numerical experiments to showcase the superiority and efficiency of the proposed PESS and LPESS preconditioners over the existing preconditioners to enhance the convergence speed of the Krylov subspace iterative process to solve three-by-three block SPPs. Our study involves a comparative analysis among GMRES process and the PGMRES process employing the proposed preconditioners ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS},$ and the existing preconditioners ${\mathscr{P}}_{BD}$, ${\mathscr{P}}_{IBD}$, ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS}$ (baseline preconditioners). The numerical results are reported from the aspect of iteration counts (abbreviated as “IT”) and elapsed CPU times in seconds (abbreviated as “CPU”). Each subsystem involving $\widehat{A}$ and $\widehat{X}$ featured in Algorithm 1 is precisely solved by applying the Cholesky factorization of the coefficient matrices. For all iterative process, the initial guess vector is ${\bf u}_{0}=0\in{\mathbb{R}}^{n+m+p}$, and the termination criterion is
@@
Before proceeding, we discuss the optimal choice for the parameters of PESS preconditioner. We can rewrite the PESS preconditioner as ${\mathscr{P}}_{PESS}=s\widetilde{{\mathscr{P}}}_{PESS},$ where $\widetilde{{\mathscr{P}}}_{PESS}=\left[\begin{matrix}\frac{1}{s}\Lambda_{1}+A&B^{T}&0\\ -B&\frac{1}{s}\Lambda_{2}&-C^{T}\\ 0&C&\frac{1}{s}\Lambda_{3}\end{matrix}\right].$ Since, the pre-factor $s$ has not much effect on the performance of PESS preconditioner, investigating the optimal parameters of ${\mathscr{P}}_{PESS}$ and $\widetilde{{\mathscr{P}}}_{PESS}$ are equivalent. As we know, a preconditioner performs efficiently when it is as close as possible to the coefficient matrix $\mathcal{A},$ observe that $\widetilde{{\mathscr{P}}}_{PESS}-\mathcal{A}={\displaystyle{\frac{1}{s}\Sigma}},$ which tends to a zero matrix as $s$ tends to positive infinity for fixed $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}.$ Thus, PESS encounters enhance efficiency for large value of $s.$ However, $s$ can not be too large as the coefficient matrix $\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B$ of the linear subsystem in step $3$ of Algorithm 1 becomes very-ill conditioned. Similar investigations also hold for LPESS preconditioner. Nevertheless, in Figure 7, we show the adaptability of the PESS and LPESS preconditioners by varying the parameter $s.$
Example 1.
We consider the three-by-three block SPP (1.1) taken from [18], where
$A=\left[\begin{matrix}I\otimes G+G\otimes I&0\\ 0&I\otimes G+G\otimes I\end{matrix}\right]\in{\mathbb{R}}^{2l^{2}\times 2l^{2}},\,\,B=\left[\begin{matrix}I\otimes F&F\otimes I\end{matrix}\right]\in{\mathbb{R}}^{l^{2}\times 2l^{2}},$ and $C=E\otimes F\in{\mathbb{R}}^{l^{2}\times l^{2}},$ where $G=\frac{1}{(l+1)^{2}}\,\mathrm{tridiag}(-1,2,-1)\in{\mathbb{R}}^{l\times l},\quad F=\frac{1}{l+1}\,\mathrm{tridiag}(0,1,-1)\in{\mathbb{R}}^{l\times l}$ , and $E=\mathrm{diag}(1,l+1,\ldots,l^{2}-l+1)\in{\mathbb{R}}^{l\times l}.$ Here, $E\otimes F$ denotes the Kronecker product of two matrices $E$ and $F,$ and $\mathrm{tridiag}(a,b,c)\in{\mathbb{R}}^{l\times l}$ denotes the $l\times l$ tridiagonal matrix with diagonal, subdiagonal, and superdiagonal entries are $a,b,$ and $c,$ respectively. For this problem, the size of the matrix $\mathcal{A}$ is $4l^{2}.$
@@
SS IT $7$ $7$ $7$ $7$ $7$
CPU 0.2780 $1.6189$ $10.6707$ $54.4493$ $184.6025$
RES $076.7967e-$ 07 $074.9738e-$ 07 $074.1383e-$ 07 $073.6189e-$ 07 $073.2556e-$ 07
RSS IT $7$ $7$ $7$ $7$ $7$
CPU $0.3248$ $1.4644$ $10.4985$ $52.4559$ $177.1895$
RES $075.2397e-$ 07 $073.6832e-$ 07 $072.7532e-$ 07 $072.1869e-$ 07 $071.8240e-$ 07
EGSS IT $5$ $5$ $4$ $4$ $4$
CPU $0.2588$ $1.3792$ $6.7546$ $34.8358$ $118.2298$
RES $086.8491e-$ 08 $071.5239e-$ 07 $079.7099e-$ 07 $076.3466e-$ 07 $074.8159e-$ 07
PESS IT $3$ $3$ $3$ $3$ $3$
CPU $0.2971$ $1.1597$ $6.6682$ $31.2670$ $103.8616$
$s=12$ RES $087.4100e-$ 08 $087.4642e-$ 08 $087.3327e-$ 08 $086.9803e-$ 08 $086.1920e-$ 08
LPESS IT $3$ $3$ $3$ $3$ $3$
CPU $0.2226$ $0.9905$ $6.1339$ $30.0975$ $98.6443$
$s=12$ RES $091.1683e-$ 09 $092.3215e-$ 09 $093.1540e-$ 09 $093.3907e-$ 09 $093.0271e-$ 09
@@
We consider the parameter choices for the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},{\mathscr{P}}_{EGSS},{\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}$ in two cases.
  • In Case I: $\alpha=0.1$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.1,\beta=1,\gamma=0.001$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=I,\Lambda_{2}=I,\Lambda_{3}=0.001I$ for ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}.$
[…]
The parameter selection in Case II is made as in [31]. The optimal value (experimentally) for $s$ in the range $[10,20]$ is determined to be $12$ for minimal CPU times. Following [18], selection of $\widehat{A}$ and $\widehat{S}$ (SPD approximations of $A$ and $S,$ respectively) for IBD preconditioner are done as follows:
$\widehat{A}=LL^{T},\quad\widehat{S}=\mathrm{diag}(B\widehat{A}^{-1}B^{T}),$
@@
Figure 1: Convergence curves for IT versus RES of PGMRES process employing BD, IBD, SS, RSS, EGSS, PESS and LPESS $(s=12)$ preconditioners in Case II for Example 1.
@@
(g) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=13$
(h) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=13$
Figure 2: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
@@
Figure 1 illustrates convergence curves pertaining to preconditioners BD, IBD, SS, RSS, EGSS, PESS and LPESS ($s=12$) for Case II with $l=16,32,48,64$ and $80$ These curves depict the relationship between the relative residue at each iteration step (RES) and IT counts.
To further illustrate the superiority of the PESS preconditioner, spectral distributions of $\mathcal{A}$ and the preconditioned matrices ${\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{RSS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for the Case II with $l=16$ and $s=13$ are displayed in Figure 2.
@@
To investigate the robustness of the proposed PESS preconditioner, we measure the condition number of the ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ which is for any nonsingular matrix $A$ defined by $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2}.$ In Figure 3, the influence of the parameter $s$ on the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ is depicted. The parameter $s$ is chosen from the interval $[5,50]$ with step size one. The data presented in Table 2 and Figures 1-3 allow us to make the following noteworthy observations:
@@
  • From Table 2, it is observed that the GMRES has a very slow convergence speed. In both Cases I and II, our proposed PESS and LPESS preconditioners outperform all other compared existing preconditioners from the aspects of IT and CPU times. For example, in Case I with $l=80,$ our proposed PESS preconditioner is $76\%$, $43\%$, $28\%$, $27\%$, and $33\%$ more efficient than the existing BD, IBD, SS, RSS, and EGSS preconditioner, respectively. Moreover, LPESS preconditioners performs approximately $78\%,$ $48\%,$ $35\%,$ $34\%,$ and $39\%$ faster than BD, IBD, SS, RSS, and EGSS preconditioner, respectively. Similar patterns are noticed for $l=16,32,48$ and $64$.
@@
  • In Figure 1, it is evident that the PESS and LPESS preconditioners have a faster convergence speed than the other baseline preconditioners when applied to the PGMRES process for all $l=16,32,48,64$ and $80$